Please help.I do not get this problem:
Graph f:x-> |x|+2 with domain D= { -2,-1,0,1,2}

Answer by Edwin McCravy(8909)


Please help.I do not get this problem:
Graph f:x-> |x| + 2 with domain D = {-2,-1,0,1,2}

All you do is substitute each of those 5 numbers in the 
domain for x in the equation f(x) = |x| + 2:

f(-2) = |-2| + 2  = 2 + 2 = 4, so plot the point (-2,4)
f(-1) = |-1| + 2  = 1 + 2 = 3, so plot the point (-1,3)
 f(0) =  |0| + 2  = 0 + 2 = 2, so plot the point  (0,2)
 f(1) =  |1| + 2  = 1 + 2 = 3, so plot the point  (1,3)
 f(2) =  |2| + 2  = 2 + 2 = 4, so plot the point  (2,4)

Here is all there is to the graph.

 

You don't connect the points since the domain just consists
of 5 numbers, and so the graph just consists of those
five points where the o's are, and no other points.

Edwin